OpenStudy (anonymous):

medal will be given and fan 7 and fifteen-sixteenths minus 4 and seven-sixteenths is equal to what number? 3 and one-sixteenth 3 and one-half 4 4 and one-half

OpenStudy (anonymous):

The equation is:\[7\frac{ 7 }{ 16 }-4\frac{ 7 }{ 16 }?\] Correct?

OpenStudy (anonymous):

I meant \[7\frac{ 15 }{ 16 }\]

OpenStudy (anonymous):

yes correct

OpenStudy (anonymous):

Okay so what you want to do is convert these mixed fractions to improper fractions. So:\[7\frac{ 15 }{ 16 }-4\frac{ 7 }{ 16 }=>\frac{ 127 }{ 16 }-\frac{ 71 }{ 16 }\]

OpenStudy (anonymous):

wrong you have to pick one the answers they gave me

OpenStudy (anonymous):

I haven't given you the answer. It's better to understand how to do it for yourself rather than not knowing how when you have a test.

OpenStudy (anonymous):

Subtract those two numbers and you get: \[\frac{ 56 }{ 16 }\]

OpenStudy (anonymous):

That gives you \[3\frac{ 1 }{ 2 }\]